# Ano : Faster is Better in Noisy Landscapes

### v1

@@ 8 Limitations and Discussion
Through the design and empirical analysis of Ano, we identified two main limitations. First, explicitly scaling updates with the instantaneous gradient norm improves robustness under noise but can also cause overshooting when transient variance spikes inflate the step size before curvature feedback stabilizes the update. We tested simple magnitude clipping, but it did not provide consistent gains. Second, learning-rate heuristics inherited from Adam do not transfer directly: because Adam’s momentum norm is typically smaller than the raw magnitude term used in Ano, the same nominal rate produces a larger effective step. In supervised vision tasks, we therefore recommend a smaller default ($10^{-}4$) instead of $10^{-}3$, which yielded more stable convergence (see Appendix B). We also explored Nesterov-style acceleration in the spirit of Adan, but initial trials suggested that combining acceleration with magnitude-aware updates increased instability rather than mitigating it. Together, these limitations highlight the need for more stable handling of the magnitude component, for example through robust norm estimation or adaptive damping, so as to prevent overshooting while retaining Ano’s resilience to noisy gradients.
@@ 9 Conclusion
We introduced Ano, an alternative to momentum-based adaptive optimizers that decouples direction and magnitude, improving robustness in noisy and non-stationary settings. Under standard smoothness and bounded-noise assumptions, we provide non-asymptotic guarantees comparable to prior analyses of sign-based methods (e.g., Signum, Lion) under similar decay schedules. Empirically, Ano delivers substantial gains in reinforcement learning and low-resource NLP tasks, while remaining competitive on low-noise tasks.
Despite these strengths, Ano can more easily overshoots due to its reliance on instantaneous magnitudes; in supervised settings we therefore recommend smaller learning rates (e.g., $10^{-4}$ rather than $10^{-3}$), as discussed in Section 8. Future work will explore adaptive damping of the magnitude term, integration with Nesterov‑style look‑ahead, and extensions to large‑scale foundation‑model pre‑training, with an eye toward dynamic hyperparameter adaptation.
@@ 10 Reproducibility Statement
The full source code (training, preprocessing, and result visualization), along with all the logs from our experiments, is available on the Github Repository  [^4] https://github.com/Adrienkgz/ano-experiments. The optimizer can also be imported by installing the library (pip install ano-optimizer) to facilitate the implementation. Data preprocessing methods are described in the relevant sections of the Experiments part (see Section 6) and are also in our source code. MuJoCo experiments were conducted on a MacBook Pro with an Apple M4 Pro chip (16-core GPU, 12-core CPU), while other experiments were run on a workstation equipped with an RTX 5090 GPU and an Intel Core Ultra 9 processor, using CUDA 12.9 and PyTorch 2.9.0. All software dependencies are specified in the requirements.txt included in the repository, ensuring exact reproducibility.
The total compute time for all experiments was approximately 2000 hours: about 125 hours for the ablation study7, 215 hours for hyperparameter tuningB, and 1700 hours for the main experiments reported in Section 6.
@@ Appendix A Anolog pseudo code
 $v_{k}=v_{k-1}-(1-\beta_{2})\cdot\text{sign}(v_{k-1}-g_{k}^{2})\cdot g_{k}^{2}$
 $x_{k+1}=x_{k}-\frac{\eta_{k}}{\sqrt{v_{k}}+\epsilon}\cdot|g_{k}|\cdot\text{sign}(m_{k})-\eta_{k}\lambda x_{k}$
Algorithm 2 Anolog without decoupled decay weights
@@ Appendix B Hyperparameter Tuning Protocol
To guarantee a balanced comparison, we carried out exhaustive grid searches on lightweight proxy tasks representative of each domain. The campaign comprised 2160 independent training runs, corresponding to 40 GPU-hours per optimiser. Adan introduces a third momentum coefficient; to keep the computational budget uniform we vary only $(\beta_{1},\beta_{3})$, which govern the first and second moment estimates, respectively, while fixing the nesterov term at its default value $\beta_{2}=0.92$. For Anolog, we tuned the learning rate and the $\beta_{2}$ value.
@@
Starting from CIFAR-10, we drew a balanced subset of 10,00010{,}000 training images and 2,0002{,}000 test images. We applied the identical augmentation pipeline used in Section 6 (random cropping, horizontal flips, and Cutout). Each hyperparameter configuration was trained for 20 epochs under five independent seeds with a ResNet-18 backbone. During training, we monitored validation accuracy and retained the single best accuracy per seed. Reported values are the mean and standard deviation of these maxima across seeds.
Figure 5: Validation-accuracy trajectories on the CIFAR-10 proxy (ResNet-18) for a representative subset of optimizers. Shaded areas denote $\pm$1 s.d. across five seeds.
For Adam, the best performance was achieved with a learning rate of $1\mathrm{e}{-3}$. The optimizer showed low sensitivity to $\beta$ values across the tested range; however, the configuration $\beta_{1}=0.9$, $\beta_{2}=0.99$ yielded the highest accuracy and lowest standard deviation across learning rates. This configuration was retained for subsequent experiments.
For Ano, optimal performance was obtained with a lower learning rate of $1\mathrm{e}{-4}$, which is atypical for ResNet training. The best results corresponded to $\beta_{1}=0.95$ and $\beta_{2}=0.995$.
For Adan, the highest accuracy was observed at $\mathrm{lr}=1\mathrm{e}{-3}$, with a strong synergy between $\beta_{1}=0.92$ and $\beta_{2}=0.9$ across all tested learning rates.
For Lion, optimal accuracy was reached with $\mathrm{lr}=1\mathrm{e}{-4}$. While performance was largely insensitive to $\beta_{1}$, $\beta_{2}=0.99$ consistently produced the best results. We therefore selected $\beta_{1}=0.9$, $\beta_{2}=0.99$.
For Grams, the selected configuration was $\mathrm{lr}=1\mathrm{e}{-3}$, $\beta_{1}=0.95$, and $\beta_{2}=0.99$ based on consistent high performance.
For Anolog, we adopted the same learning rate as for Ano ($1\mathrm{e}{-4}$), which yielded the best results. As $\beta_{2}$ had a notable impact across learning rates and values below the recommended threshold performed better, we fixed $\beta_{2}=0.95$.
@@
[…]
Figure 6: Grid-search results on MRPC used as an NLP proxy for hyper-parameter selection.
Adam achieves its highest validation performance at a learning rate of $\eta=7\times 10^{-5}$ and is only weakly sensitive to $(\beta_{1},\beta_{2})$, with the standard configuration $(0.9,0.999)$ remaining close to optimal.
Adan, in line with previous findings, favors larger learning rates, also peaking at $\eta=7\times 10^{-5}$. It displays strong sensitivity to $\beta_{1}$, with the best results obtained for $(\beta_{1},\beta_{2})=(0.95,0.92)$ and a third momentum parameter $\beta_{3}=0.96$.
Ano shows minimal sensitivity to both learning rate and momentum parameters. Its optimal configuration is found at $\eta=2\times 10^{-5}$ with $(\beta_{1},\beta_{2})=(0.9,0.999)$.
Lion reaches its best validation score at $\eta=2\times 10^{-5}$ and $\beta_{2}=0.9$. The choice of $\beta_{1}$ appears to have limited impact; we thus retain the standard $\beta_{1}=0.9$.
Grams performs best with a learning rate of $\eta=7\times 10^{-5}$ and the default $(\beta_{1},\beta_{2})=(0.9,0.999)$ configuration.
Anolog yields its highest performance at $\eta=2\times 10^{-5}$, with $\beta_{2}=0.99$ proving consistently superior. As $\beta_{1}$ had negligible influence, we fix it to 0.9.
@@
For Deep reinforcement learning, we adopted the MuJoCo HalfCheetah-v5 environment and capped training at $1\times 10^{5}$ interaction steps, a budget that suffices to reveal optimisation speed without incurring prohibitive wall-clock cost. Agents were trained with the Soft Actor–Critic (SAC) algorithm, and five seeds were run per hyperparameter setting. We report the final episodic return together with its standard deviation, and visualise learning curves in Fig. 7.
Figure 7: Learning curves on HalfCheetah-v5 using SAC. Curves show mean return; shading indicates $\pm$1 s.d. over seeds.
Figure 7 shows that while Adam appears to reach its peak performance at the relatively large step size $\eta=3\times 10^{-3}$, this result is likely overestimated due to the 100k-step evaluation window. Larger learning rates tend to accelerate early progress but stall quickly, resulting in higher variance. To ensure greater stability, we adopt the more conservative setting $\eta=3\times 10^{-4}$ with $(\beta_{1},\beta_{2})=(0.95,0.995)$, which achieves comparable performance with significantly reduced variability.
Ano exhibits low sensitivity to the tested hyperparameter range. Nevertheless, the configuration $\eta=3\times 10^{-4}$ with $(\beta_{1},\beta_{2})=(0.95,0.99)$ consistently yields the highest average score and is therefore selected.
For Adan, the best performance is obtained at $\eta=3\times 10^{-3}$ with $(\beta_{1},\beta_{2},\beta_{3})=(0.98,0.92,0.98)$, consistent with the optimizer’s preference for higher learning rates.
For Lion, the optimal configuration corresponds to $(\beta_{1},\beta_{2})=(0.92,0.99)$.
Grams performs best with a more aggressive learning rate of $\eta=3\times 10^{-3}$ and momentum coefficients $(\beta_{1},\beta_{2})=(0.9,0.95)$.
@@ C.1 Algorithmic update
For each coordinate $i\in[d]$ the optimizer maintains a first-order momentum $m_{k,i}$ and a second moment $v_{k,i}$ updated as in Yogi (Zaheer et al., 2018):
$m_{k,i}=\beta_{1,k}m_{k-1,i}+(1-\beta_{1,k})g_{k,i},\qquad v_{k,i}=v_{k-1,i}-(1-\beta_{2})\,\operatorname{sign}\!\bigl{(}v_{k-1,i}-g_{k,i}^{2}\bigr{)}\,g_{k,i}^{2},$ (1)
with $\beta_{1,k}=1-\frac{1}{\sqrt{k+1}}$ for $k\geq 1$; $\beta_{2}\in(0,1)$ and $m_{0,i}=v_{0,i}=0$. The parameter vector is then updated by
@@
All conditional expectations $\mathbb{E}_{k-1}[\,\cdot\,]=\mathbb{E}[\,\cdot\,|\mathcal{F}_{k-1}]$
@@ C.2 Standing assumptions
There exists a constant $G>0$ such that $\|\nabla f(x)\|\leq G$ for all iterates $x$.
@@
C.3 Auxiliary quantities
@@
C.4 Lemma for Probability Sign-Mismatch
@@ C.5 Bound on A
$\mathbb{E}[A_{k}]\geq\frac{\mathbb{E}[\|\nabla f(x_{k})\|^{2}_{2}]}{\sqrt{2}G+\varepsilon}-\frac{C_{v}C_{m}}{(k+1)^{1/4}},$
@@
Using the bound $0\leq v_{k-1,i}\leq 2G^{2}$, which holds under the assumptions of Zaheer et al. (2018) (see Theorem 2), we deduce:
$\mathbb{E}_{k-1}[A_{k}]\geq\frac{\|\nabla f(x_{k})\|^{2}_{2}}{\sqrt{2}G+\varepsilon}-\frac{1}{(k+1)^{1/4}}\cdot\frac{2d\sqrt{\sigma^{2}+G^{2}}\;C_{m}}{\varepsilon}$
@@
$\mathbb{E}_{k-1}[A_{k}]\geq\frac{\|\nabla f(x_{k})\|^{2}_{2}}{\sqrt{2}G+\varepsilon}-\frac{C_{v}C_{m}}{(k+1)^{1/4}},$
@@
$\mathbb{E}[A_{k}]\geq\frac{\mathbb{E}[\|\nabla f(x_{k})\|^{2}_{2}]}{\sqrt{2}G+\varepsilon}-\frac{C_{v}C_{m}}{(k+1)^{1/4}},$
@@ C.7 Main result
with $C_{sum}=C_{m}C_{v}\eta$ and $C_{LHS}=\frac{\eta}{\sqrt{2}G+\epsilon}$.
@@ D.1 Full Results on Analysis Parts
Ano $82.25\pm 0.09$ $78.78\pm 0.24$ $71.0\pm 0.33$ $65.96\pm 0.15$ $59.81\pm 0.74$
Adam $80.89\pm 0.29$ $76.08\pm 0.3$ $66.43\pm 0.08$ $60.56\pm 0.37$ $53.01\pm 0.86$
Lion $81.23\pm 0.24$ $77.69\pm 0.15$ $69.40\pm 0.11$ $64.03\pm 0.93$ $57.07\pm 0.94$
Grams $1.37\pm 0.33 77$$7.91\pm 0.03$ $70.56\pm 0.23$ $65.47\pm 0.32$ $58.8\pm 0.58$
Table 7: CIFAR-10 test accuracy (%) $\pm$ std.
D.2 Full Normalized Score for SAC
@@
Baseline
Adam $9477.97\pm 1003.20$ 87.34 $4161.08\pm 515.06$ 79.30 $\mathbf{5247.63}\pm\mathbf{105.87}$ 100 $4276.75\pm 238.68$ 83.55 $2922.12\pm 284.42$ 99.85 90.01
Adan $8796.97\pm 955.18$ 81.06 $3294.58\pm 433.50$ 62.79 $5116.14\pm 125.19$ 97.49 $3698.10\pm 572.68$ 72.24 $2509.11\pm 324.34$ 85.74 79.87
Lion $7406.13\pm 2772.84$ 68.25 $3491.82\pm 579.52$ 66.55 $85.58\pm 28.33$ 1.63 $4768.21\pm 201.84$ 93.15 $703.34\pm 657.37$ 24.03 50.72
Grams $6457.31\pm 879.37$ 59.50 $3031.85\pm 478.99$ 57.78 $5181.55\pm 105.09$ 98.74 $3386.75\pm 896.29$ 66.16 $2606.07\pm 579.75$ 89.05 74.25
Ano (Ours) $\mathbf{10851.80}\pm\mathbf{518.67}$ 100 $\mathbf{5247.02}\pm\mathbf{653.98}$ 100 $5190.86\pm 136.30$ 98.92 $\mathbf{5119.08}\pm\mathbf{420.88}$ 100 $\mathbf{2926.39}\pm\mathbf{308.72}$ 100 99.78
Anolog (Ours) $10389.50\pm 1041.29$ 95.74 $4598.67\pm 727.16$ 87.64 $5161.95\pm 268.91$ 98.37 $4748.18\pm 593.57$ 92.75 $2585.67\pm 238.48$ 88.36 92.57
@@
Adam $9874.24\pm 735.62$ 87.67 $4110.11\pm 692.96$ 76.26 $\mathbf{5308.70}\pm\mathbf{195.70}$ 100 $4369.92\pm 326.27$ 87.73 $\mathbf{3143.48}\pm\mathbf{233.52}$ 100 90.33
Adan $10133.64\pm 849.35$ 89.97 $5222.12\pm 218.59$ 96.90 $5034.40\pm 95.83$ 94.83 $\mathbf{4981.25}\pm\mathbf{184.94}$ 100 $2971.66\pm 367.72$ 94.53 95.25
Lion $7351.97\pm 3058.09$ 65.27 $2826.41\pm 1698.65$ 52.44 $86.81\pm 24.43$ 1.64 $4496.17\pm 519.54$ 90.26 $430.78\pm 260.73$ 13.70 44.66
Grams $8522.14\pm 1648.23$ 75.66 $4187.69\pm 707.71$ 77.70 $5116.82\pm 135.39$ 96.39 $4204.40\pm 836.06$ 84.40 $2870.06\pm 740.79$ 91.30 85.09
Ano (Ours) $\mathbf{11263.08}\pm\mathbf{284.18}$ 100 $\mathbf{5389.34}\pm\mathbf{353.92}$ 100 $5071.59\pm 138.17$ 95.53 $4810.42\pm 356.80$ 96.57 $2972.84\pm 418.39$ 94.57 97.34
Table 8: Comparison of the average performance (± IC95%) and normalized scores of different optimizers across MuJoCo environments.
D.3 Full Normalized Score for PPO

### v2

@@ 8 Limitations and Discussion
Through our design and empirical analysis of Ano, we identified three main limitations: First, the choice of $\beta_{2}$-decay appears particularly beneficial in reinforcement learning or highly non-stationary loss landscapes, where older gradients can be misleading and a rapid adaptation is crucial. However, in more stationary settings, such as classical supervised learning in CV and NLP, we observed that the variance estimate in vanilla Yogi often leads to more stable and effective training. Our focus on noisy, non-stationary environments motivates this design choice, though we acknowledge that its relevance to more conventional settings remains an open question. Second, by construction, Ano favors larger step sizes to improve reactivity. While this design is advantageous in non-stationary contexts, it can also introduce instability. For example, our experiments with Nesterov-style acceleration, inspired by Adan, amplified rather than mitigated this issue. Third, our experiments on classical CV and NLP tasks remain limited in scale, as Ano was primarily designed for highly non-stationary and noisy environments. In more stationary settings with longer training horizons, we observed that Adam can sometimes achieve better stability due to its smaller update steps. While these results suggest that Ano’s benefits are not restricted to DRL, assessing its relevance to large-scale CV or NLP tasks lies beyond the current scope and is left for future work.
@@ 9 Conclusion
We introduced Ano, an alternative to momentum-based adaptive optimizers that decouples direction and magnitude to improve robustness in noisy and non-stationary settings. Under standard smoothness and bounded-noise assumptions, we derive non-asymptotic guarantees comparable to existing analyses of sign-based methods (e.g., Signum, Lion) under similar decay schedules. Empirically, Ano achieves notable improvements in reinforcement learning and noisy NLP tasks while remaining competitive on low-noise benchmarks. Future work will focus on developing variance estimators tailored to supervised learning, integrating Nesterov-style look-ahead, and enhancing stability in long, stationary training regimes. [^4] LLMs (GPT, Gemini) were used for minor text editing, LaTeX formatting, informal feedback on early drafts, and retrieval of related work and references; all research ideas, analyses, and conclusions remain solely those of the authors.
@@ 10 Reproducibility Statement
All datasets used in this work are publicly available. The full source code, including training, preprocessing, and result visualization scripts, as well as all experiment logs, is released in an anonymous repository [^5] https://anonymous.4open.science/r/ano-optimizer-1645/README.md. The optimizer is also available as a pip package (PyTorch, TensorFlow, JAX) to facilitate implementation, but it isn’t include in the source code for double bind review. Data preprocessing details, hyperparameter grids, and training protocols are described in Section 6 and provided in the source code. All experiments were run with fixed random seeds on a workstation with an RTX 5090 GPU and an Intel Core Ultra 9 CPU using CUDA 12.9 and PyTorch 2.9.0.
@@ Appendix A Anolog pseudo code
 $v_{k}=\beta_{2}v_{k-1}-(1-\beta_{2})\cdot\text{sign}(v_{k-1}-g_{k}^{2})\cdot g_{k}^{2}$
 $\hat{v_{k}}=\frac{v_{k}}{1-\beta_{2}^{k}}$
 $x_{k+1}=x_{k}-\frac{\eta_{k}}{\sqrt{\hat{v_{k}}}+\epsilon}\cdot|g_{k}|\cdot\text{sign}(m_{k})-\eta_{k}\lambda x_{k}$
Algorithm 2 Anolog
@@ Appendix B Hyperparameter Tuning Protocol
To ensure a fair comparison, we conducted exhaustive grid searches over learning rates and momentum parameters on lightweight proxy tasks representative of each domain, using 5 independent seeds for each hyperparameter combination. In total, the campaign involved 2115 independent training runs: roughly 35 hours per optimizer when searching across three hyperparameters, and about 12 hours for optimizers with only two (e.g., Anolog, RMSprop). For Adan, which introduces a third momentum coefficient, we maintained a uniform computational budget by varying only $(\beta_{1},\beta_{3})$, that controlling the first and second moment estimates, while fixing the Nesterov term at its default value, $\beta_{2}=0.92$. For Anolog and RMSprop, we tuned only the learning rate and the variance term decay parameter ($\beta_{2}$ for Anolog and $\alpha$ for RMSprop). We then selected the configuration achieving the highest validation accuracy per seed, with final hyperparameters for all optimizers summarized in Table LABEL:tab:optimizers-hyperparameters.
@@
From CIFAR-10 (krizhevsky2009learning), we drew a balanced subset of 10,00010{,}000 training images and 2,0002{,}000 test images. We applied the identical augmentation pipeline used in Section 6 (random cropping, horizontal flips, and Cutout). Each hyperparameter configuration was trained for 20 epochs under five independent seeds with a ResNet-18 backbone(he2016deep).
Figure 5: Grid search on the CIFAR-10 proxy (ResNet-18) for the optimizers.
@@
We fine-tune Bert (devlin2019bert) on the MRPC benchmark(wang2019gluemultitaskbenchmarkanalysis). Although MRPC is relatively small and noisy, this characteristic amplifies the impact of optimizer hyperparameters, making it easier to reveal differences in optimization behavior that may be less pronounced on larger, more stable datasets. At the same time, its modest size keeps the experiments computationally efficient while preserving representative fine-tuning dynamics of GLUE tasks. A preliminary sweep indicated that learning rates outside the range $1!\times!10^{-5},,7!\times!10^{-5}]$ consistently led to poor accuracy; subsequent grids therefore focus on a narrower range centered at $2!\times!10^{-5}$.
Figure 6: Grid-search results on MRPC used as an NLP proxy.
@@
For Deep Reinforcement Learning, we train a SAC agent on the MuJoCo HalfCheetah-v5 environment for 100k steps, given time constraints(todorov2012mujoco; haarnoja2018soft). This setup is primarily intended to reveal the impact of different hyperparameters, especially the momentum coefficients $\beta$, though we note that the shorter horizon may favor more aggressive learning rates. To address this limitation, the main text reports the best performance between default and tuned hyperparameters for each optimizer.
Figure 7: Grid-search on HalfCheetah 100k-steps training used as a DRL proxy.
@@ C.1 Algorithmic update
For each coordinate $i\in[d]$ the optimizer maintains a first-order momentum $m_{k,i}$ and a second moment $v_{k,i}$ updated as below:
$m_{k,i}=\beta_{1,k}m_{k-1,i}+(1-\beta_{1,k})g_{k,i},\qquad v_{k,i}=\beta_{2}v_{k-1,i}-(1-\beta_{2})\,\operatorname{sign}\!\bigl(v_{k-1,i}-g_{k,i}^{2}\bigr)\,g_{k,i}^{2},$ (1)
with $\beta_{1,k}=1-\frac{1}{\sqrt{k+1}}$ for $k\geq 1$; $\beta_{2}\in[0.5,1)$ and $m_{0,i}=v_{0,i}=0$. The parameter vector is then updated by
@@
We write $\mathbb{E}_{k-1}[\cdot]:=\mathbb{E}[\cdot\mid\mathcal{F}_{k-1}]$ for the conditional expectation given the filtration $\mathcal{F}_{k-1}$.
@@ C.2 Standing assumptions
There exists a constant $G>0$ such that $|\nabla_{i}f(x_{k})|\leq G$ for all iterates $x_{k}$ and all coordinates $i$.
@@
C.3 Preliminary Lemma
Local assumption.
For this lemma only, we impose an additional assumption solely to simplify the analysis and to obtain a pointwise bound on $v_{k,i}$. Specifically, we assume $g_{k,i}\leq\tilde{G}$ for all $k,i$. This assumption is not used anywhere else in the paper and plays no role in the convergence results.
Lemma 1 (Bounds on $v_{k}$).
Fix any coordinate $i\in[d]$ and assume $v_{0}=0$ and $\beta_{2}\in[\tfrac{1}{2},1)$. Then for every $k\geq 0$,
$0\;\leq\;v_{k,i}\;\leq\;\tilde{G}^{2}.$
Proof.
The update is
$v_{k,i}=\beta_{2}v_{k-1,i}-(1-\beta_{2})\,\operatorname{sign}(v_{k-1,i}-g_{k,i}^{2})\,g_{k,i}^{2}.$
Upper bound.
If $\operatorname{sign}(v_{k-1,i}-g_{k,i}^{2})=-1$, then
$v_{k,i}=\beta_{2}v_{k-1,i}+(1-\beta_{2})g_{k,i}^{2}\leq\tilde{G}^{2}$
If $\operatorname{sign}(v_{k-1,i}-g_{k,i}^{2})=1$, then
$v_{k,i}=\beta_{2}v_{k-1,i}-(1-\beta_{2})g_{k,i}^{2}\leq\beta_{2}v_{k-1,i}\leq v_{k-1,i}.$
Starting from $v_{0}=0$, induction gives $v_{k,i}\leq G^{2}$ for all $k$.
Lower bound.
If $\operatorname{sign}(v_{k-1,i}-g_{k,i}^{2})=-1$, then $v_{k,i}\geq 0$ since it is a convex combination of nonnegative terms. If $\operatorname{sign}(v_{k-1,i}-g_{k,i}^{2})=1$, then
$v_{k,i}=\beta_{2}v_{k-1,i}-(1-\beta_{2})g_{k,i}^{2}\geq(2\beta_{2}-1)g_{k,i}^{2}\;\geq\;0$
because $\beta_{2}\geq\tfrac{1}{2}$ and $g_{k,i}^{2}\geq 0$. Thus $v_{k,i}\geq 0$ for all $k$. ∎
C.4 Auxiliary quantities
@@
C.5 Lemma for Probability Sign-Mismatch
@@ C.6 Bound on A
$\mathbb{E}[A_{k}]\geq\frac{\mathbb{E}[\|\nabla f(x_{k})\|^{2}_{2}]}{\tilde{G}+\varepsilon}-\frac{C_{v}C_{m}}{(k+1)^{1/4}},$
@@
Using lemma 1, we got $0\leq v_{k-1,i}\leq\tilde{G}^{2}$, we deduce:
$\mathbb{E}_{k-1}[A_{k}]\geq\frac{\|\nabla f(x_{k})\|^{2}_{2}}{\tilde{G}+\varepsilon}-\frac{1}{(k+1)^{1/4}}\cdot\frac{2d\sqrt{\sigma^{2}+G^{2}}\;C_{m}}{\varepsilon}$
@@
$\mathbb{E}_{k-1}[A_{k}]\geq\frac{\|\nabla f(x_{k})\|^{2}_{2}}{\tilde{G}+\varepsilon}-\frac{C_{v}C_{m}}{(k+1)^{1/4}},$
@@
$\mathbb{E}[A_{k}]\geq\frac{\mathbb{E}[\|\nabla f(x_{k})\|^{2}_{2}]}{\tilde{G}+\varepsilon}-\frac{C_{v}C_{m}}{(k+1)^{1/4}},$
@@ C.8 Main result
with $C_{sum}=C_{m}C_{v}\eta$ and $C_{LHS}=\frac{\eta}{\tilde{G}+\epsilon}$.
@@ D.1 Full Results on Analysis Parts
[…]
Table 7: CIFAR-10 test accuracy (%) with 95% confidence intervals.
@@
Default
Adam $10549.48\pm 721.55$ 97.10 $4336.64\pm 698.72$ 82.05 $5357.14\pm 211.97$ 99.29 $4462.51\pm 588.77$ 85.36 $3164.71\pm 600.48$ 89.52 90.66
RMSprop $10506.23\pm 852.19$ 96.71 $4234.37\pm 763.65$ 80.11 $5395.51\pm 126.80$ 100.00 $4160.06\pm 480.62$ 79.57 $2973.86\pm 571.05$ 84.12 88.10
Adan $7805.20\pm 1154.02$ 71.84 $2985.19\pm 1018.79$ 56.48 $5080.74\pm 305.26$ 94.17 $4092.13\pm 379.92$ 78.28 $3222.62\pm 235.25$ 91.16 78.38
Lion $9527.96\pm 805.42$ 87.70 $4948.26\pm 243.05$ 93.62 $98.22\pm 32.33$ 1.82 $4612.63\pm 367.77$ 88.23 $3087.27\pm 628.06$ 87.33 71.74
Grams $6782.60\pm 715.12$ 62.43 $3207.30\pm 531.06$ 60.68 $5104.10\pm 692.14$ 94.60 $3656.66\pm 658.82$ 69.95 $1475.34\pm 927.22$ 41.73 65.88
Ano (Ours) $10864.09\pm 1052.24$ 100.00 $5285.44\pm 729.86$ 100.00 $5255.62\pm 815.92$ 97.41 $5227.86\pm 436.49$ 100.00 $3535.32\pm 780.96$ 100.00 99.48
Anolog (Ours) $10557.05\pm 560.70$ 97.17 $5089.12\pm 522.94$ 96.29 $5242.78\pm 173.98$ 97.17 $4606.02\pm 478.36$ 88.11 $3314.12\pm 539.95$ 93.74 94.50
@@
Adam $8243.01\pm 2750.47$ 69.70 $5050.53\pm 471.12$ 90.82 $5224.24\pm 339.87$ 100.00 $4429.62\pm 668.97$ 90.83 $2968.40\pm 696.52$ 82.62 86.79
RMSprop $10096.62\pm 2379.00$ 85.37 $3509.99\pm 827.41$ 63.12 $64.97\pm 42.44$ 1.24 $4583.19\pm 969.52$ 93.98 $2031.80\pm 771.10$ 56.55 60.05
Adan $10822.40\pm 475.75$ 91.51 $5239.69\pm 270.96$ 94.22 $4792.62\pm 904.44$ 91.74 $4686.83\pm 502.28$ 96.11 $3514.42\pm 143.57$ 97.82 94.28
Lion $10482.06\pm 1018.86$ 88.63 $4848.41\pm 821.79$ 87.18 $1349.15\pm 1322.56$ 25.82 $4876.76\pm 253.22$ 100.00 $3592.87\pm 70.26$ 100.00 80.33
Grams $10533.70\pm 866.69$ 89.07 $4607.59\pm 505.08$ 82.85 $5147.04\pm 487.55$ 98.52 $4644.45\pm 498.08$ 95.24 $3147.82\pm 605.03$ 87.61 90.66
Ano (Ours) $11826.22\pm 700.46$ 100.00 $5561.17\pm 400.26$ 100.00 $5158.30\pm 313.97$ 98.74 $4804.34\pm 359.02$ 98.51 $3226.05\pm 504.58$ 89.79 97.41
Anolog (Ours) $11198.28\pm 771.94$ 94.69 $5095.64\pm 722.28$ 91.63 $3137.03\pm 1335.43$ 60.05 $4563.86\pm 834.12$ 93.58 $3321.91\pm 472.93$ 92.46 86.48
Best Version
Adam $10549.48\pm 721.55$ 97.10 $4336.64\pm 698.72$ 82.05 $5357.14\pm 211.97$ 99.29 $4462.51\pm 588.77$ 85.36 $3164.71\pm 600.48$ 88.08 90.38
RMSprop $10506.23\pm 852.19$ 96.71 $4234.37\pm 763.65$ 80.11 $5395.51\pm 126.80$ 100.00 $4160.06\pm 480.62$ 79.57 $2973.86\pm 571.05$ 82.77 87.83
Adan $10822.40\pm 475.75$ 99.62 $5239.69\pm 270.96$ 99.13 $4792.62\pm 904.44$ 88.83 $4686.83\pm 502.28$ 89.65 $3514.42\pm 143.57$ 97.82 95.01
Lion $10482.06\pm 1018.86$ 96.48 $4848.41\pm 821.79$ 91.73 $1349.15\pm 1322.56$ 25.01 $4876.76\pm 253.22$ 93.28 $3592.87\pm 70.26$ 100.00 81.30
Grams $10533.70\pm 866.69$ 96.96 $4607.59\pm 505.08$ 87.18 $5147.04\pm 487.55$ 95.39 $4644.45\pm 498.08$ 88.84 $3147.82\pm 605.03$ 87.61 91.20
Ano (Ours) $10864.09\pm 1052.24$ 100.00 $5285.44\pm 729.86$ 100.00 $5255.62\pm 815.92$ 97.41 $5227.86\pm 436.49$ 97.41 $3535.32\pm 780.96$ 98.40 99.16
Anolog (Ours) $10557.05\pm 560.70$ 97.17 $5089.12\pm 522.94$ 96.29 $5242.78\pm 173.98$ 97.17 $4606.02\pm 478.36$ 97.17 $3314.12\pm 539.95$ 92.24 94.20
Table 8: Comparison of the average performance ($\pm$ CI95%) and normalized scores of different optimizers across MuJoCo environments.
D.2 Full Normalized Score for PPO